# Error Correction in Dynamical Codes

### v1

@@ 4.5 Proof for the Distance Algorithm
If the update is according to Case 1 for $C_{1},V_{1}$, it must also be updated by Case 1 for $C_{2},V_{2}$. $m$ is added to $V_{i}$, so $\langle V_{1}\rangle=\langle V_{2}\rangle$. It is clear that elements in $\langle C_{1}\rangle\cap\langle V_{1}\rangle$ still differ from $\langle C_{2}\rangle\cap\langle V_{2}\rangle$ up to $\langle\{v_{1},\cdots,v_{k}\}\rangle$ and are equivalent in finding the unmasked stabilizers. If it is Case 2 for $C_{1},V_{1}$, then it must be Case 2 for $C_{2},V_{2}$. We add $\{v_{1},\cdots,v_{k}\}$ to both $C_{1}$ and $C_{2}$, so that they generate the same group, and update them using the stabilizer update rules in Lemma 3, then the sets are updated to $\{v_{1},\cdots,v_{k}\}\cup C_{i}^{\prime}\cup\{m\}$, with $C_{i}\to C_{i}^{\prime}$, $i\in\{1,2\}$. The updated sets still generate the same group as the stabilizer update rules are basis independent. Removing $\{v_{1},\cdots,v_{k}\}\cup\{m\}$ from both sets, it is clear that elements in $C_{i}^{\prime}$ are equivalent up to elements in $\langle\{v_{1},\cdots,v_{k}\}\rangle$. $C_{i}^{\prime}$ are exactly the updated $C_{i}$ where $\{v_{1},\cdots,v_{k}\}$ has not been added to the sets. Since $\langle V_{1}\rangle=\langle V_{2}\rangle$ and $\{v_{1},\cdots,v_{k}\}\in\langle V_{i}\rangle$ after the update, an element in $\langle C_{1}\rangle$ still differs from $\langle C_{2}\rangle$ by elements in $\langle\{v_{1},\cdots,v_{k}\}\rangle$.
@@
Therefore, after the update, for an element $c_{1}$ in $\langle C_{1}\rangle\cap\langle V_{1}\rangle$ that has an associated stabilizer generator $s$, we can still find $c_{2}\in\langle C_{2}\rangle$ with the same associated stabilizer $s$, and since $c_{1}$ and $c_{2}$ differ by an element $v\in\langle V\rangle$ before the update, we still have $v\in V$ after the update, so $c_{2}\in\langle C_{2}\rangle\cap\langle V_{2}\rangle$ after the update. If it is Case 3 for both $C_{1},V_{1}$ and $C_{2},V_{2}$, the updates only change $V$. Since the update for $V$ is the same as the stabilizer update rules with $V$ as the set of stabilizers, this update is independent of the choice of basis for $V$, so after the update, $\langle V_{1}\rangle=\langle V_{2}\rangle$. For any element $c_{1}$ in $\langle C_{1}\rangle$ with associated stabilizer generator $s_{c}$, $c_{1}\cdot v_{c_{1}}\in\langle C_{2}\rangle$ for some $v_{c_{1}}\in\langle v_{1},\cdots,v_{k}\rangle$ with the same associated stabilizer generator. Then this implies $[m,v_{c_{1}}]=0$, otherwise, $m$ anti-commutes with some elements in either $C_{1}$ or $C_{2}$. Therefore, $\{v_{1},\cdots,v_{k}\}$ remains in $\langle V_{i}\rangle$, so it follows that the statement remains true for Case 3. If it is Case 4 for $C_{1},V_{1}$ and either Case 3 or 4 for $C_{2},V_{2}$, by the same reasoning as in Case 3, after the update by $m$, $\langle V_{1}\rangle=\langle V_{2}\rangle$. Suppose an element $a\in V_{1}$ is removed and an element $b\in V_{2}$ is removed. By the stabilizer update rules, some elements in $C_{1}$ are updated with $a$ and some in $C_{2}$ are updated with $b$, so the group generated by the associated stabilizer generators remains the same for both $C_{i}$. The update works the same for $C_{1},C_{2}$ if we consider updating $G_{1}=\{a,b,v_{1},\cdots,v_{k}\}\cup C_{1}$ and $G_{2}=\{b,a,v_{1},\cdots,v_{k}\}\cup C_{2}$ as the stabilizer sets. Updating by the measurement $m$ such that $a$ is removed from $G_{1}$ and $b$ is removed from $G_{2}$, we have $G_{1}\to G_{1}^{\prime}=\{m,ab,v_{1}^{\prime},v_{2}^{\prime},\cdots,v_{k}^{\prime}\}\cup C_{1}^{\prime}$ and $G_{2}\to G_{2}^{\prime}=\{m,ab,\tilde{v}_{1},\tilde{v}_{2},\cdots,\tilde{v}_{k}\}\cup C_{2}^{\prime}$, where $v_{i}\in G_{1}$ is updated to $v_{i}^{\prime}$ and $v_{i}\in G_{2}$ is updated to $\tilde{v}_{i}$. First note that $G_{1},G_{2}$ generate the same group, so after the update, $\langle G_{1}^{\prime}\rangle=\langle G_{2}^{\prime}\rangle$. Further, the updated $C_{i}$ is the same as $C_{i}^{\prime}$, as both are updated by the same elements $a$ or $b$. Since $\{b,a,v_{1},\cdots,v_{k}\}\subset\langle V\rangle$, after the update, both $\{m,ab,v_{1}^{\prime},v_{2}^{\prime},\cdots,v_{k}^{\prime}\}$ and $\{m,ab,\tilde{v}_{1},\tilde{v}_{2},\cdots,\tilde{v}_{k}\}$ are subsets of the updated $\langle V\rangle$. If we remove $\{m,ab,v_{1}^{\prime},v_{2}^{\prime},\cdots,v_{k}^{\prime}\}$ and $\{m,ab,\tilde{v}_{1},\tilde{v}_{2},\cdots,\tilde{v}_{k}\}$ from $G_{1}^{\prime}$ and $G_{2}^{\prime}$ respectively, then the elements in $C_{1}^{\prime}$ differ from those in $C_{2}^{\prime}$ by some elements in $\{ab\}\cup\{v_{1}^{\prime},v_{2}^{\prime},\cdots,v_{k}^{\prime}\}\cup\{\tilde{v}_{1},\tilde{v}_{2},\cdots,\tilde{v}_{k}\}$ which is a subset of the updated $\langle V\rangle$. Therefore, every element in the updated $C_{1}$ can still be expressed as elements in the updated $C_{2}$ and elements in the updated $\langle V\rangle$, and vice versa.
@@
[…]
@@
If the measurements consist of only those that fall under Case 1, then the statement holds trivially. Let $\{m_{1},m_{2},\cdots,m_{q}\}$ be a series of measurements. Suppose we were able to obtain the outcome for the stabilizer associated with $u\in\langle C\rangle\cap\langle V\rangle$ after measuring $m_{i}$ with $1\leq i\leq q$. Let $s_{u}\in\langle S_{0}\rangle$ be the unmasked stabilizer associated with $u$. If the measurements end here, calculating $\langle C\rangle\cap\langle V\rangle$ gives the syndrome for $s_{u}$. Using Lemma 43, the stabilizers for $C$ and $V$ are changed so that $u$ is in the new sets of $C$ and $V$.
@@
By induction, suppose all the unmasked stabilizers that can be obtained in the first $k$ measurements can be found using the distance algorithm in Subsection 4.3. We want to show that this also holds for the first $k+1$ measurements. Suppose $s_{0}\in S_{0}$ is a stabilizer generator whose syndrome can be obtained using the first $k+1$ measurements, so $O(s_{0})$ is given by the outcome of some of the first $k$ measurements and $O(m_{k+1})$, where $O(m_{i})\in{\pm 1}$ is the outcome of the measurement $m_{i}$. Since $m_{k+1}$ must be measured last, the outcome of $m_{k+1}$ must be deterministic. We know that $s_{0}$ is in the stabilizer set in round 0, and that its outcome is pre-determined, if no error occurs. This means that we know in the situation of no error on the code what the outcome of $s_{0}$ will be if we measure it at round 0. But because errors may have occurred, the rounds of measurements allow us to compare the two outcomes for $s_{0}$, to determine the errors that have occurred at round 0. Thus, the outcome of $m_{k+1}$ must be deterministic in nature if we have fixed the outcomes for $m_{i}$ with $i\leq k$ and errors $E$ on the code. In other words, $m_{k+1}$ is in $\langle C\cup V\rangle$, prior to the $(k+1)^{\mathrm{th}}$ measurements.
@@
The outcome for $c_{k+1}$ can be obtained from measuring $m_{k+1}$. $c_{k+1}$ must not be identity operator, otherwise $O(s_{0})$ can be obtained from the first $k$ measurements. Let $s_{k+1}$ be the stabilizer associated with $c_{k+1}$, then $s_{k+1}\in\langle S_{0}\rangle$ is an unmasked stabilizer generator. Since $c_{k+1}=v_{k+1}\cdot m_{k+1}$, $c_{k+1}\in\langle V\rangle$ too. If $s_{k+1}=s_{0}$, then $O(s_{0})$ is found. For the case where $s_{k+1}\neq s_{0}$, we first note that the outcomes of $s_{k+1}$ and $s_{0}$ are independent of each other, so to find $s_{0}$ using $s_{k+1}$, the only way is that $s_{0}\cdot s_{k+1}$ has a known outcome using the first $k$ measurements. Since $s_{0}\cdot s_{k+1}$ is unmasked in the first $k$ measurements, by the induction hypothesis, $s_{0}\cdot s_{k+1}$ can be found using the distance algorithm after some measurements. By Lemma 44, we know that if a stabilizer is unmasked by computing $\langle C\rangle\cap\langle V\rangle$ at some point, then it can be obtained by computing $\langle C\rangle\cap\langle V\rangle$ after all future measurements. Therefore, there exists some element $\tilde{c}$ in $\langle C\rangle$, such that it has the associated stabilizer $s_{0}\cdot s_{k+1}$, and $\tilde{c}$ is in $\langle C\rangle\cap\langle V\rangle$ after $k+1$ measurements. Then $\tilde{c}\cdot c_{k+1}$ is an element in $\langle C\rangle\cap\langle V\rangle$ with the stabilizer $s_{0}$ associated with it. Thus, we showed by induction that $s_{0}$ can be unmasked using the distance algorithm, if it is an unmasked stabilizer. ∎
@@
Suppose $c\in C$ is removed after the $k^{\mathrm{th}}$ measurement and $s_{c}$ associated with $c$ has not been unmasked by the distance algorithm. Since Theorem 42 and Theorem 45 demonstrate that every unmasked stabilizer from the first $k$ measurements can be obtained using the distance algorithm, $s_{c}$ is either temporarily or permanently masked by the measurement sequence. Applying Lemma 44, after the $k^{\mathrm{th}}$ measurement, the set of stabilizers that can be unmasked by the distance algorithm are those associated with the generators of $\langle C\rangle\cap\langle V\rangle$. If the measurement sequence ends here, then any generator that has not been unmasked but is associated with some stabilizer in $\langle C\rangle$ is temporarily masked. This is because we can measure and unmask it in the future, for example, by directly measuring the generator in the next round. Further, using the update rules in the distance algorithm, any stabilizer generator that is not associated with an element in $\langle C\rangle$ at this point will still not be associated with any element after any additional sequence of measurements. Thus, it will not be unmasked by the distance algorithm. Applying Theorem 45 again, this implies that the generator cannot be unmasked by all future measurements and therefore is permanently masked. Since $s_{c}$ is not associated with any generator in $\langle C\rangle$ after the $k^{\mathrm{th}}$ measurement, it is permanently masked. From here, we can infer that the set of stabilizers from $\langle S_{0}\rangle$ that we can unmask from the given sequence and further measurements can only be the set $\langle T\rangle\langle U\rangle$. Therefore, the permanently masked stabilizers $P$ is given by $\langle S_{0}\rangle\backslash\langle T\rangle\langle U\rangle$. This precisely corresponds to the stabilizers that are removed from $C$ due to Case 2 in Step 2 of the distance algorithm with associated stabilizers that have not been unmasked yet.
@@ 5 Initialization of a Floquet Code
We prove this by induction. We first show that this is true for the first two cycles of measurements. It holds that $S_{1}^{(1)}$ is a subset of $S_{2}^{(1)}$, since $S_{1}^{(1)}$ consists of a single measurement $m_{1}$, and after measuring $m_{1}$ in the second cycle, $m_{1}$ must be in $S_{2}^{(1)}$. To show that after each measurement, the updated $S_{1}^{(1)}$ will still be a subset of the updated $S_{2}^{(1)}$, we apply step 2 of the Distance Algorithm with $C$ = $S_{2}^{(1)}\backslash S_{1}^{(1)}$ and $V=S_{1}^{(1)}$, with the measurement sequence given by $M\backslash\{m_{1}\}$. After the $i^{\mathrm{th}}$ measurement, $V$ is updated to $S_{1}^{(i)}$. Thus, $S_{1}^{(i)}\subset V^{(i)}\cup C^{(i)}=S_{2}^{(i)}$, where $V^{(i)}$ denotes the updated $V$ and $C^{(i)}$ denotes the updated $C$, after the $i^{\mathrm{th}}$ measurement. Suppose this is true for the first $k$ cycles, we want to show that it is true for the first $k+1$ cycles. By the induction hypothesis, $S_{j-1}^{(i)}$ is a subset of $S_{j}^{(i)}$, $\forall i\leq l$, $\forall j<k-1$. We want to show that $S_{k}^{(i)}\subseteq S_{k+1}^{(i)},\forall i$. Since $S_{k-1}^{(l)}$ is a subset of $S_{k}^{(l)}$, we can run step 2 of the distance algorithm to two sets of $C,V$, initializing the first set as $C_{a_{0}}=\{\},V_{a_{0}}=S_{k-1}^{l}$, and the second set as $C_{b_{0}}=S_{k}^{(l)}\backslash S_{k-1}^{(l)},V_{b_{0}}=S_{k-1}^{(l)}$. We have $V_{a_{0}}=V_{b_{0}}$, but $C_{a_{0}}\subseteq C_{b_{0}}$ and $C_{b_{0}}\cup V_{b_{0}}=S_{k}^{(l)}$. Running the distance algorithm with the set of measurements $M$, it is easy to see that $V_{a_{0}}=V_{b_{0}}$ after all $i$ measurements, and since $C_{a_{0}}$ remains empty, $V_{a_{0}}\cup C_{a_{0}}\subseteq V_{b_{0}}\cup C_{b_{0}}$. Therefore, $S_{k}^{(i)}\subseteq S_{k+1}^{(i)},\forall i$. ∎
@@
The condition for the set of stabilizer generators to increase with a measurement $m$ is $[m,s]=0,\forall s\in S$ and $m\notin\langle S\rangle$. Since $S_{j}^{(i)}\subseteq S_{j+1}^{(i)}$, if the condition holds for $S_{j+1}^{(i)}$, then it must also hold for $S_{j}^{(i)}$, thus an increase in the number of stabilizer generators can only happen due to the set of measurements that causes the number of generators to increase in the previous cycle. ∎
Theorem 48 implies that in the worst case scenario, it takes $n-1$ cycles to initialize a Floquet code with $n$ stabilizer generators.
@@
There exists a measurement sequence for a Floquet code that takes $n-1$ cycles to fully initialize for a code with $n$ stabilizer generators.
@@
We show this by giving a recursive construction of measurement sequence such that it takes $n-1$ cycles to fully initialize the code with $n$ stabilizers.
Let $\{s_{i}:1\leq i\leq n\}$ denote a set of independent stabilizer generators for the fully initialized code. Let $d_{s_{i}}$ denote the destabilizer for $s_{i}$. The construction works in a way that $s_{i}$ is added to the set of stabilizer generators after $i-1$ cycles and each cycle increases the number of generators by 1 even though the same set of measurements is made in each cycle. The rest of the generators are updated in a way that $s_{1}$ is mapped to $s_{1}$ and $s_{i}$ is mapped to $s_{i+1}$ after one cycle. The mapping is simply a way to keep track of how the stabilizer generators change with measurements as shown in Figure 6 for illustration.

### v2

@@ 5.5 Proof for the Distance Algorithm
If the update is according to Case 1 for $C_{1},V_{1}$, it must also be updated by Case 1 for $C_{2},V_{2}$. $m$ is added to $V_{i}$, so $\langle V_{1}\rangle=\langle V_{2}\rangle$. It is clear that elements in $\langle C_{1}\rangle\cap\langle V_{1}\rangle$ still differ from $\langle C_{2}\rangle\cap\langle V_{2}\rangle$ up to $\langle\{v_{1},\cdots,v_{k}\}\rangle$ and are equivalent in finding the unmasked stabilizers.
If it is Case 2 for $C_{1},V_{1}$, then it must be Case 2 for $C_{2},V_{2}$. We add $\{v_{1},\cdots,v_{k}\}$ to both $C_{1}$ and $C_{2}$, so that they generate the same group, and update them using the stabilizer update rules in Lemma 3, then the sets are updated to $\{v_{1},\cdots,v_{k}\}\cup C_{i}^{\prime}\cup\{m\}$, with $C_{i}\to C_{i}^{\prime}$, $i\in\{1,2\}$. The updated sets still generate the same group as the stabilizer update rules are basis independent. Removing $\{v_{1},\cdots,v_{k}\}\cup\{m\}$ from both sets, it is clear that elements in $C_{i}^{\prime}$ are equivalent up to elements in $\langle\{v_{1},\cdots,v_{k}\}\rangle$. $C_{i}^{\prime}$ are exactly the updated $C_{i}$ where $\{v_{1},\cdots,v_{k}\}$ has not been added to the sets. Since $\langle V_{1}\rangle=\langle V_{2}\rangle$ and $\{v_{1},\cdots,v_{k}\}\in\langle V_{i}\rangle$ after the update, an element in $\langle C_{1}\rangle$ still differs from $\langle C_{2}\rangle$ by elements in $\langle\{v_{1},\cdots,v_{k}\}\rangle$.
@@
Therefore, after the update, for an element $c_{1}$ in $\langle C_{1}\rangle\cap\langle V_{1}\rangle$ that has an associated stabilizer generator $s$, we can still find $c_{2}\in\langle C_{2}\rangle$ with the same associated stabilizer $s$, and since $c_{1}$ and $c_{2}$ differ by an element $v\in\langle V\rangle$ before the update, we still have $v\in V$ after the update, so $c_{2}\in\langle C_{2}\rangle\cap\langle V_{2}\rangle$ after the update.
@@
If it is Case 3 for both $C_{1},V_{1}$ and $C_{2},V_{2}$, the updates only change $V$. Since the update for $V$ is the same as the stabilizer update rules with $V$ as the set of stabilizers, this update is independent of the choice of basis for $V$, so after the update, $\langle V_{1}\rangle=\langle V_{2}\rangle$. For any element $c_{1}$ in $\langle C_{1}\rangle$ with associated stabilizer generator $s_{c}$, $c_{1}\cdot v_{c_{1}}\in\langle C_{2}\rangle$ for some $v_{c_{1}}\in\langle v_{1},\cdots,v_{k}\rangle$ with the same associated stabilizer generator. Then this implies $m,v_{c_{1}}]=0$, otherwise, $m$ anti-commutes with some elements in either $C_{1}$ or $C_{2}$. Therefore, $\{v_{1},\cdots,v_{k}\}$ remains in $\langle V_{i}\rangle$, so it follows that the statement remains true for Case 3.
If it is Case 4 for $C_{1},V_{1}$ and either Case 3 or 4 for $C_{2},V_{2}$, by the same reasoning as in Case 3, after the update by $m$, $\langle V_{1}\rangle=\langle V_{2}\rangle$. Suppose an element $a\in V_{1}$ is removed and an element $b\in V_{2}$ is removed. By the stabilizer update rules, some elements in $C_{1}$ are updated with $a$ and some in $C_{2}$ are updated with $b$, so the group generated by the associated stabilizer generators remains the same for both $C_{i}$. The update works the same for $C_{1},C_{2}$ if we consider updating $G_{1}=\{a,b,v_{1},\cdots,v_{k}\}\cup C_{1}$ and $G_{2}=\{b,a,v_{1},\cdots,v_{k}\}\cup C_{2}$ as the stabilizer sets. Updating by the measurement $m$ such that $a$ is removed from $G_{1}$ and $b$ is removed from $G_{2}$, we have $G_{1}\to G_{1}^{\prime}=\{m,ab,v_{1}^{\prime},v_{2}^{\prime},\cdots,v_{k}^{\prime}\}\cup C_{1}^{\prime}$ and $G_{2}\to G_{2}^{\prime}=\{m,ab,\tilde{v}_{1},\tilde{v}_{2},\cdots,\tilde{v}_{k}\}\cup C_{2}^{\prime}$, where $v_{i}\in G_{1}$ is updated to $v_{i}^{\prime}$ and $v_{i}\in G_{2}$ is updated to $\tilde{v}_{i}$.
First note that $G_{1},G_{2}$ generate the same group, so after the update, $\langle G_{1}^{\prime}\rangle=\langle G_{2}^{\prime}\rangle$. Further, the updated $C_{i}$ is the same as $C_{i}^{\prime}$, as both are updated by the same elements $a$ or $b$. Since $\{b,a,v_{1},\cdots,v_{k}\}\subset\langle V\rangle$, after the update, both $\{m,ab,v_{1}^{\prime},v_{2}^{\prime},\cdots,v_{k}^{\prime}\}$ and $\{m,ab,\tilde{v}_{1},\tilde{v}_{2},\cdots,\tilde{v}_{k}\}$ are subsets of the updated $\langle V\rangle$.
If we remove $\{m,ab,v_{1}^{\prime},v_{2}^{\prime},\cdots,v_{k}^{\prime}\}$ and $\{m,ab,\tilde{v}_{1},\tilde{v}_{2},\cdots,\tilde{v}_{k}\}$ from $G_{1}^{\prime}$ and $G_{2}^{\prime}$ respectively, then the elements in $C_{1}^{\prime}$ differ from those in $C_{2}^{\prime}$ by some elements in $\{ab\}\cup\{v_{1}^{\prime},v_{2}^{\prime},\cdots,v_{k}^{\prime}\}\cup\{\tilde{v}_{1},\tilde{v}_{2},\cdots,\tilde{v}_{k}\}$ which is a subset of the updated $\langle V\rangle$. Therefore, every element in the updated $C_{1}$ can still be expressed as elements in the updated $C_{2}$ and elements in the updated $\langle V\rangle$, and vice versa.
@@
In the second case, we instead obtain $\alpha=\beta+1(\mathrm{mod}\>2)$, but again $c_{1}^{\prime}\cdot c_{2}^{\prime}=v\cdot a^{\alpha}\cdot b^{\alpha+1}\in\langle V\rangle$, since both $v\cdot b$ or $v\cdot a$ is in $\langle V\rangle$ and $ab\in\langle V\rangle$.
@@
If the measurements consist of only those that fall under Case 1, then the statement holds trivially.
Let $\{m_{1},m_{2},\cdots,m_{q}\}$ be a series of measurements. Suppose we were able to obtain the outcome for the stabilizer associated with $u\in\langle C\rangle\cap\langle V\rangle$ after measuring $m_{i}$ with $1\leq i\leq q$. Let $s_{u}\in\langle S_{0}\rangle$ be the unmasked stabilizer associated with $u$. If the measurements end here, calculating $\langle C\rangle\cap\langle V\rangle$ gives the syndrome for $s_{u}$.
Using Lemma 43, the stabilizers for $C$ and $V$ are changed so that $u$ is in the new sets of $C$ and $V$.
@@
By induction, suppose all the unmasked stabilizers that can be obtained in the first $k$ measurements can be found using the distance algorithm in Subsection 5.3. We want to show that this also holds for the first $k+1$ measurements.
Suppose $s_{0}\in S_{0}$ is a stabilizer generator whose syndrome can be obtained using the first $k+1$ measurements, so $O(s_{0})$ is given by the outcome of some of the first $k$ measurements and $O(m_{k+1})$, where $O(m_{i})\in{\pm 1}$ is the outcome of the measurement $m_{i}$.
Since $m_{k+1}$ must be measured last, the outcome of $m_{k+1}$ must be deterministic. We know that $s_{0}$ is in the stabilizer set in round 0, and that its outcome is pre-determined, if no error occurs. This means that we know in the situation of no error on the code what the outcome of $s_{0}$ will be if we measure it at round 0. But because errors may have occurred, the rounds of measurements allow us to compare the two outcomes for $s_{0}$, to determine the errors that have occurred at round 0. Thus, the outcome of $m_{k+1}$ must be deterministic in nature if we have fixed the outcomes for $m_{i}$ with $i\leq k$ and errors $E$ on the code.
In other words, $m_{k+1}$ is in $\langle C\cup V\rangle$, prior to the $(k+1)^{\mathrm{th}}$ measurements.
@@
The outcome for $c_{k+1}$ can be obtained from measuring $m_{k+1}$. $c_{k+1}$ must not be identity operator, otherwise $O(s_{0})$ can be obtained from the first $k$ measurements. Let $s_{k+1}$ be the stabilizer associated with $c_{k+1}$, then $s_{k+1}\in\langle S_{0}\rangle$ is an unmasked stabilizer generator. Since $c_{k+1}=v_{k+1}\cdot m_{k+1}$, $c_{k+1}\in\langle V\rangle$ too.
If $s_{k+1}=s_{0}$, then $O(s_{0})$ is found. For the case where $s_{k+1}\neq s_{0}$, we first note that the outcomes of $s_{k+1}$ and $s_{0}$ are independent of each other, so to find $s_{0}$ using $s_{k+1}$, the only way is that $s_{0}\cdot s_{k+1}$ has a known outcome using the first $k$ measurements.
Since $s_{0}\cdot s_{k+1}$ is unmasked in the first $k$ measurements, by the induction hypothesis, $s_{0}\cdot s_{k+1}$ can be found using the distance algorithm after some measurements. By Lemma 44, we know that if a stabilizer is unmasked by computing $\langle C\rangle\cap\langle V\rangle$ at some point, then it can be obtained by computing $\langle C\rangle\cap\langle V\rangle$ after all future measurements. Therefore, there exists some element $\tilde{c}$ in $\langle C\rangle$, such that it has the associated stabilizer $s_{0}\cdot s_{k+1}$, and $\tilde{c}$ is in $\langle C\rangle\cap\langle V\rangle$ after $k+1$ measurements.
Then $\tilde{c}\cdot c_{k+1}$ is an element in $\langle C\rangle\cap\langle V\rangle$ with the stabilizer $s_{0}$ associated with it. Thus, we showed by induction that $s_{0}$ can be unmasked using the distance algorithm, if it is an unmasked stabilizer.
@@
Suppose $c\in C$ is removed after the $k^{\mathrm{th}}$ measurement and $s_{c}$ associated with $c$ has not been unmasked by the distance algorithm.
Since Theorem 42 and Theorem 45 demonstrate that every unmasked stabilizer from the first $k$ measurements can be obtained using the distance algorithm, $s_{c}$ is either temporarily or permanently masked by the measurement sequence.
Applying Lemma 44, after the $k^{\mathrm{th}}$ measurement, the set of stabilizers that can be unmasked by the distance algorithm are those associated with the generators of $\langle C\rangle\cap\langle V\rangle$. If the measurement sequence ends here, then any generator that has not been unmasked but is associated with some stabilizer in $\langle C\rangle$ is temporarily masked. This is because we can measure and unmask it in the future, for example, by directly measuring the generator in the next round. Further, using the update rules in the distance algorithm, any stabilizer generator that is not associated with an element in $\langle C\rangle$ at this point will still not be associated with any element after any additional sequence of measurements. Thus, it will not be unmasked by the distance algorithm. Applying Theorem 45 again, this implies that the generator cannot be unmasked by all future measurements and therefore is permanently masked.
Since $s_{c}$ is not associated with any generator in $\langle C\rangle$ after the $k^{\mathrm{th}}$ measurement, it is permanently masked.
[…]
@@ 6 Initialization of a Floquet Code
In this section, we do not consider the effects of errors on initialization. We instead assume that the error-correcting properties of the Floquet code to initialize and maintain the code space. Instead, we focus on the question of how the code space evolves under the sequence of measurements defining the code and how long it takes to reach a periodically repeating state.
@@
We prove this by induction.
[…]
Suppose this is true for the first $k$ cycles, we want to show that it is true for the first $k+1$ cycles. By the induction hypothesis, $S_{j-1}^{(i)}$ is a subset of $S_{j}^{(i)}$, $\forall i\leq l$, $\forall j<k-1$. We want to show that $S_{k}^{(i)}\subseteq S_{k+1}^{(i)},\forall i$. Since $S_{k-1}^{(l)}$ is a subset of $S_{k}^{(l)}$, we can run step 2 of the distance algorithm to two sets of $C,V$, initializing the first set as $C_{a_{0}}=\{\},V_{a_{0}}=S_{k-1}^{l}$, and the second set as $C_{b_{0}}=S_{k}^{(l)}\backslash S_{k-1}^{(l)},V_{b_{0}}=S_{k-1}^{(l)}$. We have $V_{a_{0}}=V_{b_{0}}$, but $C_{a_{0}}\subseteq C_{b_{0}}$ and $C_{b_{0}}\cup V_{b_{0}}=S_{k}^{(l)}$. Running the distance algorithm with the set of measurements $M$, it is easy to see that $V_{a_{0}}=V_{b_{0}}$ after all $i$ measurements, and since $C_{a_{0}}$ remains empty, $V_{a_{0}}\cup C_{a_{0}}\subseteq V_{b_{0}}\cup C_{b_{0}}$. Therefore, $S_{k}^{(i)}\subseteq S_{k+1}^{(i)},\forall i$.
@@
The condition for the set of stabilizer generators to increase with a measurement $m$ is $m,s]=0,\forall s\in S$ and $m\notin\langle S\rangle$. Since $S_{j}^{(i)}\subseteq S_{j+1}^{(i)}$, if the condition holds for $S_{j+1}^{(i)}$, then it must also hold for $S_{j}^{(i)}$, thus an increase in the number of stabilizer generators can only happen due to the set of measurements that causes the number of generators to increase in the previous cycle. ∎
Theorem 48 implies that in the worst case scenario, it takes n1n-1 cycles to initialize a Floquet code with $n$ stabilizer generators.
@@
There exists a measurement sequence for a Floquet code that takes n1n-1 cycles to fully initialize for a code with $n$ stabilizer generators.
@@
We show this by giving a recursive construction of measurement sequence such that it takes n1n-1 cycles to fully initialize the code with $n$ stabilizers.
Let $\{s_{i}:1\leq i\leq n\}$ denote a set of independent stabilizer generators for the fully initialized code. Let $d_{s_{i}}$ denote the destabilizer for $s_{i}$. The construction works in a way that $s_{i}$ is added to the set of stabilizer generators after i1i-1 cycles and each cycle increases the number of generators by 1 even though the same set of measurements is made in each cycle. The rest of the generators are updated in a way that $s_{1}$ is mapped to $s_{1}$ and $s_{i}$ is mapped to $s_{i+1}$ after one cycle. The mapping is simply a way to keep track of how the stabilizer generators change with measurements as shown in Figure 6 for illustration.